# How do Large Language Models Handle Multilingualism?

### v2

@@ 3.4 Verify the Reasoning Structure in MWork
Deactivating Method
Table 3 shows the result of the reasoning task, where we deactivate $6$ sets of neurons. We adhere to the previous logic of selecting deactivation settings, with the exception that we do not conduct an independent experiment on deactivating neurons in the understanding layer, as its functionality has already been verified. Details are listed in Appendix G.
@@ 3.5 Verify the Knowledge Extraction Structure in MWork
Table 4 shows the result of the knowledge question answering task, where we deactivate $5$ sets of neurons. Similarly, we exclude the deactivation of neurons in layers that have already been verified and instead concentrate on the self-attention structure and feed-forward structure in the task-solving layer. Details are listed in Appendix G.
@@ 3.6 Verify the Generation Structure in MWork
Table 5 shows the result of the generation task, where we deactivate $3$ sets of neurons. Since all previous layers have been verified, we solely deactivate neurons in the generation layer and compare them with randomly selected neurons. Details are listed in Appendix G.
@@
4 Multilingual Enhancement with MWork
We employ Mistral-7b-v0.1 for enhancement to eliminate the interference of instruction fine-tuning, and select causal language modeling as our training task. We create a dataset comprising $\{100,200,400,800\}$ randomly selected documents for each language, extracted from the Wikipedia corpus. [^4] https://dumps.wikimedia.org/ Figure 4 shows the results of enhancement on high-resource languages (De, Fr, Zh, Es, Ru). The numbers represent the sizes of the training corpus when fine-tuning language-specific neurons, while "Random" represents the fine-tuning of an equivalent number of randomly chosen neurons using a corpus of $400$. Our findings reveal that fine-tuning with a few hundred documents yields significant performance improvements on multilingual tasks: $3.4\%$ on MGSM, $4.4\%$ on XQuAD, $4.3\%$ on X-CSQA, and $2.3\%$ on XLSum. Moreover, English performance is enhanced by an average of $3.7\%$ across all tasks. These results further confirm the effectiveness of MWork in interpreting structure functionality for LLM’s multilingual query handling, offering precise and independent methods for multilingual enhancement. When fine-tuning with 800 documents, the performance deteriorates compared to using 400 documents. This drop can be attributed to the incorporation of additional knowledge, which disrupts the original knowledge distribution and leads to overfitting of the model to Wikipedia. This can be addressed by mixing data from more sources such as textbooks or websites.
In addition, we verify the effectiveness of such enhancement method on low-resource languages, given that low-resource performance is relatively low with the original model. We select four languages including Vietnamese (Vi), Thai (Th), Arabic (Ar), and Swahili (Sw), covering languages with both latin and non-latin scripts and having corresponding testing set in our considered benchmarks. The model was then evaluated on four benchmarks, and the result shown in Table 4 is the average among tasks. It is evident that the fine-tuning method using language-specific neurons enhances the model’s multilingual performance in low-resource languages by an average of $2.2\%$. Notably, the improvement of $3.5\%$ in English performance is observed even without an English training corpus, indicating the effectiveness of the distinct language responsibilities assigned to neurons.
@@ 4 Multilingual Enhancement with MWork
Table 6: Enhancement is achieved by fine-tuning Mistral-7b-v0.1 model utilizing $400$ documents from each language correspondingly. The results are averaged across four tasks. Performance on English (“En”) is obtained by averaging the results from three fine-tuned models.
@@ 5 Related Work
In the era of Language and Linguistic Models (LLMs), numerous studies have been conducted to develop multilingual benchmarks (Zhang et al., 2023a), enhance multilingual performance without parameter adjustments through translation (Liang et al., 2023; Huang et al., 2023), aligning representations (Nguyen et al., 2023a; Salesky et al., 2023), prompting (Li et al., 2023b; Tanwar et al., 2023). Furthermore, certain works focus on improving multilingual abilities for a single task via cross-lingual transfer (Kim et al., 2017; Lin et al., 2019; Pfeiffer et al., 2020; Zhao et al., 2024b), while others aim to enhance multilingual proficiency by continuous training in one language to obtain mono-lingual LLMs (Cui et al., 2023a), or in multiple domain languages to obtain domain-lingual LLMs (Nguyen et al., 2023b). Additionally, some works achieve multilingual LLMs by training from scratch (Muennighoff et al., 2023). However, these studies are limited to specific task types or require substantial training corpora due to a lack of comprehensive understanding of the multilingual mechanisms of LLMs.
Conventional interpretability research investigates the significance of input features with their corresponding outputs (Vig, 2019; Hewitt and Liang, 2019; Qiu et al., 2020). In the era of LLMs, one brunch of work includes efforts to understand knowledge storage, with (Geva et al., 2021) initiating the study of the feed-forward layer as a knowledge base. Subsequent work has furthered this by altering neuron values (Dai et al., 2022b), mapping embeddings to words (Geva et al., 2022), modifying inputs to recover embeddings (Meng et al., 2022), and analyzing attention heads (Li et al., 2023a). Another line of research centers on the self-attention layer, examining its connection to reasoning capability (Hou et al., 2023; Stolfo et al., 2023; Friedman et al., 2023) by contrasting the reasoning tree based on attention weights.
@@ 6 Conclusion
In this work, we examine how LLMs handle multilingualism. The proposed multilingual workflow (MWork) suggests that LLMs initially understand queries by converting multilingual inputs into English, think in English in intermediate layers while incorporating multilingual knowledge, and generate responses aligned with the original language in the final layers. The validity of MWork is verified using Parallel Language-specific Neuron Detection (PLND), which identifies activated neurons for different languages without labeled data. By detecting language-specific neurons and fine-tuning them with a small training corpus, MWork enhances multilingual abilities in specific languages without compromising others, resulting in significant average improvements across tasks.
Limitation and Impact Statements
Our experiments are mainly conducted on models with a size of approximately 7 billion parameters (i.e., vicuna-7b-v1.5 and mistral-7b-v1.0), primarily due to computational resource constraints. Testing our methods on larger models could potentially yield additional insights, particularly in understanding the scalability of our proposed framework. Furthermore, our exploration into enhancing the multilingual abilities of language models through our framework was preliminary. Expanding our experiments to include a broader array of languages, especially those considered low-resource, could better demonstrate the effectiveness of our proposed framework. Moreover, extending the scope of our experiments to evaluate other capabilities such as reasoning and multilingual knowledge extraction with specific datasets could provide a more comprehensive picture of the potential benefits of our approach.
Our paper has the potential to significantly enhance the multilingual ability of LLMs. By effectively improving their performance across all languages, it can enable the development of real multilingual models that excel in various applications, promoting better communication and understanding among diverse linguistic communities.
@@ Appendix C Interrelation of Language-Specific Neurons Across Languages
We can observe that in both Mistral and Vicuna, the intersection with English from other languages is relatively limited, suggesting that English possesses a predominant number of language-specific neurons. Additionally, there is a pronounced tendency for languages belonging to the same family to demonstrate a higher degree of overlap with each other, such as Spanish, French, and English. Moreover, the feed-forward structure typically exhibits a higher degree of consistency in overlap across various languages, due to the shared world knowledge embedded within the neurons that is accessible to multiple languages.
Appendix D Prompts
@@
Appendix E Original Performance
@@
Appendix F Hyper-parameters
@@
Appendix G Detailed Experiment Results
G.1 Detailed Experiment Settings
@@ G.2 Detailed Result
Appendix H Analysis on Different Multilingual LLMs
To further verify MWork, we examine two more types of multilingual LLMs, including BLOOMZ (Muennighoff et al., 2022), a hyper-multilingual LLM claiming to support 46 languages, and Chinese Llama (Cui et al., 2023b), a bilingual LLM focusing on English and Chinese.
Hyper-Multilingual LLMs
From Figure 1(b) we observe that BLOOMZ also tends to default to English as its “thinking” language. Therfore, it can support multiple languages mainly because it has the ability to understand and generate these languages while using English as the language of thinking. Additionally, Figure 7 illustrates the degree of neuron overlap among languages within both the self-attention and feed-forward structures of BLOOMZ. In contrast to the findings shown in Figure 5, there is a marked reduction in overlap, indicating that individual languages maintain a higher degree of independence and do not extensively share neurons with one another.
Figure 6: Overlapping ratio of language-specific neurons in BLOOMZ
Bilingual LLMs
We employ Chinese Llama (Cui et al., 2023b), which extends existing vocabulary and incorporate secondary pre-training using Chinese data and fine-tune the model with Chinese instruction datasets. However, this intensive training can lead to a degradation in performance for languages other than Chinese. As depicted in Figure 7, Chinese predominates as the primary language for thinking processing and knowledge extraction across all languages. Consequently, the absence of language-specific neurons results in the transformation of it into a Chinese-centric LLM.

### v3

@@ 3.4 Verify the Reasoning Structure in MWork
Deactivating Method
Table 3 shows the result of the reasoning task, where we deactivate $6$ sets of neurons. We adhere to the previous logic of selecting deactivation settings, with the exception that we do not conduct an independent experiment on deactivating neurons in the understanding layer, as its functionality has already been verified. Details are listed in Appendix I.
@@ 3.5 Verify the Knowledge Extraction Structure in MWork
Table 4 shows the result of the knowledge question answering task, where we deactivate $5$ sets of neurons. Similarly, we exclude the deactivation of neurons in layers that have already been verified and instead concentrate on the self-attention structure and feed-forward structure in the task-solving layer. Details are listed in Appendix I.
@@ 3.6 Verify the Generation Structure in MWork
Table 5 shows the result of the generation task, where we deactivate $3$ sets of neurons. Since all previous layers have been verified, we solely deactivate neurons in the generation layer and compare them with randomly selected neurons. Details are listed in Appendix I.
@@
Table 6: Enhancement is achieved by fine-tuning Mistral-7b-v0.1 model utilizing $400$ documents from each language correspondingly. The results are averaged across four tasks. Performance on English (“En”) is obtained by averaging the results from four fine-tuned models.
@@
4 Multilingual Enhancement with MWork
We employ Mistral-7b-v0.1 for enhancement to eliminate the interference of instruction fine-tuning, and select causal language modeling as our training task. We create a dataset comprising $\{100,200,400,800\}$ randomly selected documents for each language, extracted from the Wikipedia corpus (Foundation, ). Figure 4 shows the results of enhancement on high-resource languages (De, Fr, Zh, Es, Ru). The numbers represent the sizes of the training corpus when fine-tuning language-specific neurons, while "Random" represents the fine-tuning of an equivalent number of randomly chosen neurons using a corpus of $400$. Our findings reveal that fine-tuning with a few hundred documents yields significant performance improvements on multilingual tasks: $3.4\%$ on MGSM, $4.4\%$ on XQuAD, $4.3\%$ on X-CSQA, and $2.3\%$ on XLSum. Moreover, English performance is enhanced by an average of $3.7\%$ across all tasks. These results further confirm the effectiveness of MWork in interpreting structure functionality for LLM’s multilingual query handling, offering precise and independent methods for multilingual enhancement. When fine-tuning with 800 documents, the performance deteriorates compared to using 400 documents. This drop can be attributed to the incorporation of additional knowledge, which disrupts the original knowledge distribution and leads to overfitting of the model to Wikipedia. This can be addressed by mixing data from more sources such as textbooks or websites.
In addition, we verify the effectiveness of such enhancement method on low-resource languages, given that low-resource performance is relatively low with the original model. We select four languages including Vietnamese (Vi), Thai (Th), Arabic (Ar), and Swahili (Sw), covering languages with both latin and non-latin scripts and having corresponding testing set in our considered benchmarks. The model was then evaluated on four benchmarks, and the result shown in Table 3.6 is the average scores among tasks. It is evident that the fine-tuning method using language-specific neurons enhances the model’s multilingual performance in low-resource languages by an average of $2.2\%$. Notably, the improvement of $3.5\%$ in English performance is observed even without an English training corpus, indicating the effectiveness of the distinct language responsibilities assigned to neurons.
@@ 5 Related Work
In the era of LLMs, numerous studies have been conducted to develop multilingual benchmarks (Zhang et al., 2023a), enhance multilingual performance without parameter adjustments through translation (Liang et al., 2023; Huang et al., 2023), aligning representations (Nguyen et al., 2023a; Salesky et al., 2023), prompting (Li et al., 2023b; Tanwar et al., 2023). Furthermore, certain works focus on improving multilingual abilities for a single task via cross-lingual transfer (Kim et al., 2017; Lin et al., 2019; Pfeiffer et al., 2020; Zhao et al., 2024b), while others aim to enhance multilingual proficiency by continuous training in one language to obtain mono-lingual LLMs (Cui et al., 2023), or in multiple domain languages to obtain domain-lingual LLMs (Nguyen et al., 2023b). Additionally, some works achieve multilingual LLMs by training from scratch (Muennighoff et al., 2023). However, these studies are limited to specific task types or require substantial training corpora due to a lack of comprehensive understanding of the multilingual mechanisms of LLMs.
Conventional interpretability research investigates the significance of input features with their corresponding outputs (Vig, 2019; Hewitt and Liang, 2019; Qiu et al., 2020). In the era of LLMs, one brunch of work includes efforts to understand knowledge storage, with (Geva et al., 2021) initiating the study of the feed-forward layer as a knowledge base. Subsequent work has furthered this by altering neuron values (Dai et al., 2022), mapping embeddings to words (Geva et al., 2022), modifying inputs to recover embeddings (Meng et al., 2022), and analyzing attention heads (Li et al., 2023a). Another line of research centers on the self-attention layer, examining its connection to reasoning capability (Hou et al., 2023; Stolfo et al., 2023; Friedman et al., 2023) by contrasting the reasoning tree based on attention weights.
@@ 6 Conclusion
In this work, we examine how LLMs handle multilingualism. The proposed multilingual workflow (MWork) suggests that LLMs initially understand queries by converting multilingual inputs into English, reason in English in intermediate layers while incorporating multilingual knowledge, and generate responses aligned with the original language in the final layers. The validity of MWork is verified using Parallel Language-specific Neuron Detection (PLND), which identifies activated neurons for different languages without labeled data. By detecting language-specific neurons and fine-tuning them with a small training corpus, MWork enhances multilingual abilities in specific languages without compromising others, resulting in significant improvements across tasks.
Acknowledgement
This work was substantially supported by DAMO Academy through DAMO Academy Research Intern Program. This research is partially supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG2-TC-2023-010-SGIL) and the Singapore Ministry of Education Academic Research Fund Tier 1 (Award No: T1 251RES2207).
@@ Appendix C Interrelation of Language-Specific Neurons Across Languages
We can observe that in both Mistral and Vicuna, the intersection with English from other languages is relatively limited (i.e., the first row of each figure), suggesting that English possesses a predominant number of language-specific neurons. Additionally, there is a pronounced tendency for languages belonging to the same family to demonstrate a higher degree of overlap with each other, such as Spanish, French, and English.
Appendix D Analysis on Different Multilingual LLMs
We further examine two more types of multilingual LLMs, including BLOOMZ (Muennighoff et al., 2023), a hyper-multilingual LLM claiming to support 46 languages, and Chinese Llama (Cui et al., 2023), a bilingual LLM focusing on English and Chinese.
Hyper-Multilingual LLMs
Figure 7 illustrates the degree of neuron overlap among languages within both the self-attention and feed-forward structures of BLOOMZ. In contrast to the findings shown in Figure 5, there is a marked reduction in overlap, indicating that individual languages maintain a higher degree of independence and do not extensively share neurons with one another.
Figure 6: Overlapping ratio of language-specific neurons in BLOOMZ
Bilingual LLMs
We employ Chinese Llama (Cui et al., 2023), which extends existing vocabulary and incorporate secondary pre-training using Chinese data and fine-tune the model with Chinese instruction datasets. However, this intensive training can lead to a degradation in performance for languages other than Chinese. As depicted in Figure 7, Chinese predominates as the primary language for reasoning processing and knowledge extraction across all languages. Consequently, the absence of language-specific neurons results in the transformation of it into a Chinese-centric LLM.
Appendix E Language-Agnostic Neurons
We initially implement a radical deactivation approach, wherein we specifically deactivate overlapping elements between each language and English. These elements precisely correspond to the intersecting neurons in the first column of Figure 5. Presented below are the comprehensive findings pertaining to Mistral. Our evaluation is centered around the reasoning task, which is recognized as the most indicative and challenging assessment for the model. We compare under the optimal “deactivating” method, which involves deactivating all language-specific neurons except those in S-ATTN.
Table 8: Performance of deactivating language-specific neurons without overlapped between English.
Language Eng non-Eng $\Delta_{\text{Eng}}$ $\Delta_{\text{non-Eng}}$ $\Delta\uparrow$
All language-specific neurons $46.2$ $18.3$ $+0.2$ $-8.0$ $+8.2$
LSN without overlapped between English $45.8$ $20.2$ $-0.2$ $-6.1$ $+5.9$
As evident by Table 8, the performance of English remains stable, contrasting sharply with the significant decline in the performance of multilingual. Removing overlapped neurons, as opposed to deactivating all language-specific neurons, leads to a less pronounced drop, yet the impact remains noteworthy. This demonstrates that overlapped neurons are not language-agnostic; they are not utilized for general comprehension and logical reasoning. Otherwise, the fundamental reasoning capacity and performance in multilingual contexts would remain unaffected. In addition, we retained the language-specific neurons that overlapped in all languages, meaning that we removed them from the language-specific neurons to be deactivated. Detailed results follow.
Table 9: Performance of deactivating language-specific neurons without all languages overlapped.
Language Eng non-Eng $\Delta_{\text{Eng}}$ $\Delta_{\text{non-Eng}}$ $\Delta\uparrow$
All language-specific neurons $46.2$ $18.3$ $+0.2$ $-8.0$ $+8.2$
LSN without all languages overlapped $45.6$ $18.7$ $-0.4$ $-7.6$ $+7.2$
The neurons that overlap across all languages only account for $0.02\%$ of the total number of neurons. From the results in Table 9, we can see that the performance is almost the same as deactivating all language-specific neurons. This further proves that these neurons are not language-agnostic neurons, but only a subset of language-specific neurons.
Appendix F Prompts
@@
Appendix G Original Performance
@@
Appendix H Hyper-parameters
@@
Appendix I Detailed Experiment Results
I.1 Detailed Experiment Settings